OpenStudy (anonymous):

3x + 5 – 2x – 4 ------ ------ = 3 6 5

OpenStudy (anonymous):

\[(3x+5)/6\]is \[(3/6)x +(5/6)\] you can figure out the rest.

OpenStudy (anonymous):

ahhh i'm confused :/

OpenStudy (anonymous):

first, u have to diminish the fraction such that: \[1\div x + 1 \div y = t\] this equal to \[1\div x \times xy + 1\div y \times xy = t \times xy\] thus \[y + x = txy\]